OpenStudy (anonymous):

Find the type and number of solutions for the equation: 2x2 + 18 + 12x = 0

OpenStudy (theviper):

\[2x^2+18+12x=0\] It is Quadratic Polynomial.

OpenStudy (anonymous):

2x^2 + 18 + 12x = 0 2(x^2+9+6x)=0 x^2+9+6x=0 (x+3)^2=0 x+3=0 x=-3

OpenStudy (theviper):

& it has two solutions :)

OpenStudy (anonymous):

\[x ^{2}+6x+9=0\]

OpenStudy (theviper):

got it @tonio5ilva ?

OpenStudy (anonymous):

Now either: factor, complete the square or use the quadratic equation to solve. gl

OpenStudy (anonymous):

There is only 1 solution.

OpenStudy (theviper):

if D=0 then yes it has 1 solution :)